OpenStudy (anonymous):

Terrence normally uses 7/8 teaspoon of dried basil to make his own salad dressing. He decides to change the recipe to use less dried basil. He decides that the flavor is perfect when using a total of 3/4 teaspoon. Write a subtraction equation that could be used to determine the amount of basil, b, Terrence excluded. Solve for b. Describe your steps.

OpenStudy (anonymous):

3/4-7/8=b 2/2*(3/4)-7/8=b You need to get a common denominator in order to subtract them (6/8)-(7/8)=b Now it's simple numerator math: 1/8=b